# GenKnowSub: Improving Modularity and Reusability of LLMs through General Knowledge Subtraction

### v1

@@ Abstract
Large language models (LLMs) often struggle with zero-shot generalization, and several modular approaches have been proposed to address this challenge. Yet, we hypothesize that a key limitation remains: the entanglement of general knowledge and task-specific adaptations. To overcome this, we propose a modular framework that disentangles these components by constructing a library of task-specific LoRA modules alongside a general-domain LoRA. By subtracting this general knowledge component from each task-specific module, we obtain residual modules that focus more exclusively on task-relevant information—a method we call general knowledge subtraction (GenKnowSub). Leveraging the refined task-specific modules and the Arrow routing algorithm (Ostapenko et al., 2024), we dynamically select and combine modules for new inputs without additional training. Our studies on the Phi-3 model and standard Arrow as baselines reveal that using general knowledge LoRAs derived from diverse languages, including English, French, and German, yields consistent performance gains in both monolingual and cross-lingual settings across a wide set of benchmarks. Further experiments on Phi-2 demonstrate how GenKnowSub generalizes to weaker LLMs. The complete code and data are available at https://github.com/saharsamr/Modular-LLM.
@@
Mohammadtaha Bagherifard2  Sahar Rajabi1thanks:   Equal contribution.  Ali Edalat [^1] footnotemark: 1  Yadollah Yaghoobzadeh1,2 1Electrical and Computer Engineering Department, University of Tehran, Tehran, Iran 2Tehran Institute for Advanced Studies, Khatam University, Tehran, Iran taha.bagheri98@gmail.com, {sahar.rajabi, ali.edalat, y.yaghoobzadeh}@ut.ac.ir
@@ Appendix B Experiments on Phi-2
Method Setting PIQA BOOLQ SWAG HSWAG ARC-E ARC-C WG OQA BBH Avg
Phi-2 78.99 81.16 63.50 66.75 82.11 53.51 56.51 44.00 48.00 63.84
Arrow 79.65 81.13 65.75 66.41 83.38 54.84 60.85 48.60 54.75 65.15
GenKnowSub $En$ 79.97 80.12 65.58 66.75 84.38 54.51 61.24 49.80 54.00 66.26
$Avg$ 80.47 78.47 66.10 67.96 84.03 56.19 60.69 47.80 54.00 66.19
Table 3: Comparison of accuracy across different methods using Phi-2 as the base model on English datasets in a zero-shot setting, with Accuracy as the reported metric.
Method Setting HSWAG ARC-C XNLI MMLU Avg
German Phi-2 28.78 23.84 34.50 24.19 27.83
Arrow 28.75 24.83 32.50 26.91 28.25
GenKnowSub $En$ 28.33 24.55 33.33 24.91 27.78
$Avg$ 28.42 23.49 34.33 25.50 27.93
French Phi-2 33.33 26.84 34.16 24.77 29.77
Arrow 32.91 27.51 34.16 25.68 30.06
GenKnowSub $En$ 33.50 25.50 31.83 26.11 29.23
$Avg$ 32.25 24.16 37.50 25.00 29.73
Table 4: Performance comparison of different methods using Phi-2 as the base model on multilingual datasets in a zero-shot setting for German and French, with Accuracy as the reported metric.

### v2

@@ Abstract
Large language models often struggle with zero-shot generalization, and several modular approaches have been proposed to address this challenge. Yet, we hypothesize that a key limitation remains: the entanglement of general knowledge and task-specific adaptations. To overcome this, we propose a modular framework that disentangles these components by constructing a library of task-specific LoRA modules alongside a general-domain LoRA. By subtracting this general knowledge component from each task-specific module, we obtain residual modules that focus more exclusively on task-relevant information—a method we call general knowledge subtraction (GenKnowSub). Leveraging the refined task-specific modules and the Arrow routing algorithm (Ostapenko et al., 2024), we dynamically select and combine modules for new inputs without additional training. Our studies on the Phi-3 model and standard Arrow as baselines reveal that using general knowledge LoRAs derived from diverse languages, including English, French, and German, yields consistent performance gains in both monolingual and cross-lingual settings across a wide set of benchmarks. Further experiments on Phi-2 demonstrate how GenKnowSub generalizes to weaker LLMs. The complete code and data are available at https://github.com/saharsamr/Modular-LLM.
@@
Mohammadtaha Bagherifard2  Sahar Rajabi1thanks:   Equal contribution.  Ali Edalat [^1] footnotemark: 1 Yadollah Yaghoobzadeh1,2 1School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran 2Tehran Institute for Advanced Studies, Khatam University, Tehran, Iran taha.bagheri98@gmail.com, {sahar.rajabi, ali.edalat, y.yaghoobzadeh}@ut.ac.ir